# Learning to Infer Generative Template Programs for Visual Concepts

### v1

@@ C.1 Omniglot
$UBlock\xrightarrow{}\texttt{UNION}(SBlock,SBlock)\mid SBlock;$
$SBlock\xrightarrow{}(SymBlock\mid CBlock\mid MBlock\mid SBlock\mid PBlock);UBlock$
@@
$SBlock\xrightarrow{}\texttt{Scale}(wt,wf,ht,hf)$
@@
$ShapeBlock\xrightarrow{}PBlock;ShapeBlock\mid\texttt{FILL}~{}\mid~{}\texttt{END}$
@@
This DSL creates shape structures by defining cuboids, and arranging them through attachment. Cuboids are instantiated with the Cuboid command. Each Attach command moves one command to connect to previous part, indicated by $cube_{n}$ at location specified by the other parameters of the command. This language supports the creation of reflectional symmetry groups (Reflect) and translational symmetry groups Translate. Of note, we allow the DSL to expand hierarchically, so that Cuboids can become the bounding volume of their own sub-program (represented above with the return to the START block). These nested sub-programs are allowed to be set to a completely filled mode (FILL) or instead expand into empty space if immediately followed by the END operator. Differing from other languages, we only allow HOLE tokens to replace these START tokens that define sub-program structures, to better match the hierarchical processes by which manufactured shapes are commonly modeled.
@@ D.1 Architecture Details
When we consume a primitive soup of input, we use a different architecture based on a Transformer Encoder (Vaswani et al., 2017). We assume that each primitive is a cuboid with 6 dimensions that describe its 3D position and size. We linearize these primitive attributes, and lift each of them to dimension 16 with a 2-layer MLP. Following this we add a learned positional encoding to each attribute based on its attribute type. We then have another ‘positional encoding’ that is produced by concatenating all of the attributes of each primitive (in the lifted dimension) and sending this feature through a 2-layer MLP that outputs an embedding of the same size as the lifted dimension, which then gets summed back into each attribute. This scheme allows us to avoid worrying about how the primitives are ordered, while still allowing the attention scheme of the network to differentiate which attributes belong to which primitives. We send this tokenized representation through a standard Transformer encoder network, where we prepend the sequence with four ‘dummy’ tokens. Each token attends to every other token, and we treat the representations output in the indices of the four ‘dummy’ tokens as the visual tokenization. These dummy tokens build up a representation that attends of the entire input in much the same way as [CLS] tokens have been employed. Note that this encoder assumes a maximum number of primitives as input, which we set to 20. If the input scene does not have 20 primitives, we leave these entries as zeros, and then don’t attend over those corresponding positions in the sequence.
@@ Appendix E Training Details
We implement our networks in PyTorch (Paszke et al., 2017). We run all experiments on a NVIDIA GeForce RTX 3090 with 24GB of GPU memory, and 64 GB of RAM. During pretraining we set the batch size to max out GPU memory, this amounts to sizes of 32 for the 2D layout domain, 40 for Omniglot domain, 32 for the shape domain with a primitive soup input and 16 for the shape domain with voxel inputs (of size $64^{3}$). Note that this batch size is effectively multiplied by 5 for the ExpansionNet and ParamNet as we train on visual input groups of size 5. During fine-tuning we set the beam size to 20 for all methods, except for the shape-voxels variant, which we set to 10 to avoid maxing out VRAM.
@@
We finetune our inference networks with the procedure described in Section 3.3. For each concept in the training set, we sample a group of visual inputs (at random) from the concept, and record our inference results to produce the LEST and ST dataset. In this way if there are K concepts in $X^{*}$, the size of the ST and LEST data on each training step will also be K. Differing from this, in the wake-sleep step of our finetuning procedure we can generate an arbitrarily large number of paired data by sampling our generative model. We find that sampling a large number of ‘dreams’ is helpful for our finetuning procedure, so we set the number of example $TP$ to sample in each training step to 30000. This typically takes between 1 and 2 hours, different slightly for each domain. To encourage the ‘dreams’ we sample to cover a wide-distribution, we design a negative rejection step where we resample any ‘dream’ that either creates an already generated $TP$ or $X^{G}$. We find this rejection criteria is triggered at relatively infrequent rates ($\sim$5of the time).
Once we’ve created the ST, LEST and WS datasets, we use them to finetune our inference networks with cross entropy loss. We train over this datasets for multiple ‘epochs’, where every 5th epoch we run our updated inference networks over concepts from the validation set. We use the Objective $O$ from this validation inference to decide when to break out of the training step, and return to the inference step. This early stopping inference procedure always back tracks to the version of the inference network that achieved the best $O$ measure in a given training step. We use a patience of 10 epochs, and finetune for at most 50 epochs.
Overall, we run our finetuning procedure to convergence for 25, 17, 32 inference-training loops for the layout, omniglot and shape domains respectively. This corresponded with 565, 450, 620 inference training ‘epochs’ for these domains. For the weights of our objective function $O$, we normalize each reconstruction metric to values typically between 0 and 1, and then we set $\lambda_{1}$ to 1.0 and $\lambda_{2}$ to 0.001. Moreover, when calculating the divergent description length between each Template Program and its respective program instantiations, we discard counting any parameter-type for which we don’t support parameter relations. For instance, as we don’t allow float variables to use parametric relations (see Section B.2), we do not penalize these variables under $O$, because the $TP$ has no opportunity to constrain them.
@@ F.1 Few-shot Generation
Some of these networks directly compare the generated samples to reference set for each concept. Frechet distance (FD) (Heusel et al., 2017) measures the distributional similiarity between two distributions of enodings. Minimum Matching Distance (MMD) measures the average minimum distance of each member of the reference set to any member of the generated set. Coverage (Cov) measures the percentage of reference set members who are the nearest neighbors to at least one member of the generated set.
We calculate all of the above metrics with respect to a latent space that is domain-specific. Specifically, for each domain, we train a visual auto-encoder to learn how to reconstruct ‘random’ scenes from that domain. For the layout domain these are randomly placed primitives. For omniglot, these are randomly placed strokes. For shapes, these are randomly place cuboids primitives. We train each of these networks to convergence on 500000 random scenes with a small bottleneck layer size (e.g. 100).
For the layout and omniglot domain we train simple classifier networks to learn a K-way classification over all of the concepts present in the domain. For Omniglot, we train on 19 examples from each of the 1623 characters in the dataset, and keep 1 example from each concept as a validation set. Our classifier achieves a 82.4% validation accuracy after convergence. For layout, we train over 95 examples from each concepts in a 20-way classification task over meta-concepts; we reach 99.9% validation accuracy on a held out set of 5 examples per concept. The class confidence metric (Conf) is then computed by taking each generated output, running it through the classifier, and then recording the probability that the classifier predicts for the index of the input concept. Note that this metric is not dependant on the reference set of examples.
@@ F.1.1 Perceptual Study
Participants were either shown 50 Omniglot character comparisons or 25 shape comparisons. We visualized shape comparisons with a simple rendering style of the primitive outputs produced by each method (for time considerations). Our supplemental material includes all of the images that were used in our perceptual study.
@@ F.2.1 Ground-Truth Segmentations
We make use of Omniglot stroke data to produce the ground-truth segmentations for characters. We treat each stroke pattern broken by ‘BREAK’ annotations as a separate segment (Lake et al., 2019a). Then, as humans vary in the ways that they order strokes to draw characters, for each test set character we run a clustering procedure to try to find valid and consistent segmentation groupings. We first filter for finding groups of characters that use the same number of strokes, and more than a single stroke (otherwise the co-segmentation task is trivial). Then we encode each stroke with a 4 dimensional feature: its length, its angle, its starting x position, and its starting y position. We run an unsupervised clustering algorithm over this feature representation (Ester et al., 1996), identify if there is any cluster with more than 5 character members, and then take 5 characters from this cluster as a co-segmentation task (where our feature-wise distance creates a correspondence across the strokes of this group). This automatic process generates 306 co-segmentation tasks from the 659 concept in the Omniglot test set. We manually inspect the generated tasks, and filter out 22 cases where our clustering identified a group that did not have consistent stroke expression. This leaves us with 284 cosegmentation tasks that we use in our experiments. All of these tasks are visualized in the supplemental material
@@ F.2.2 Group Parsing
Template Programs support parsing by inferring instantiations from a shared $TP$ that explain a group of visual inputs. As each instantiated program $z$ uses the function call structure specified by a Template Program, we can find correspondences in the visual outputs. We create a corresponding group for each primitive type that the Template Program defines: these are created by the PRIM command for the layout domain, the DRAW command for Omniglot and the Cuboid command for 3D Shapes. Note that HOLE tokens are always treated as a construct that creates primitive types. Any command that operates over this primitive type will inherent their corresponding part index (e.g. symmetry operations), exclusing combinators like Union.
@@
BAE-NET creates corresponding group parses by performing an argmax over the last layer of an implicit network that is trained to solve occupancy tasks. This implicit network can be run over any spatial position, and assign this input point to one of it part ‘slots’.
@@ G.2 GNS (Feinman & Lake, 2021)
Using the provided code as reference, we re-implemented the system with a simple CNN architecture, following the VAE framing as described in (Rybkin et al., 2020). Though we spent a fair amount of time tuning hyper-parameters, as evidenced by the quantitative results in Table 1 we were unable to achieve competitive performance.
@@
Once we have trained this VQ-VAE for each domain, we can learn our arVHE model. Like the VHE model, and our system, it learns by sampling random visual groups from the same concept. Following the procedure described in the VHE paper and code, we encode these visual concepts with a visual encoder, take a mean embedding, then use this embedding to condition an auto-regressive generation process, where a sequence of VQ-VAE tokens are predicted (corresponding to another input example from the same concept). We train this network with cross-entropy loss, on the discretized VQ-VAE tokens. For an apples-to-apples comparison against our method, the arVHE baseline uses the same visual encoders that our method uses (Section D). For predicting 3D shapes as a sequence of primitives, we instead just task the VQ-VAE model with predicting discretrized versions of each primitive attribute, where the primitives are randomly ordered (this allows us to skip the VQ-VAE step in this setting).
@@
We follow the author’s implemented released at: https://github.com/czq142857/BAE-NET. We take their architecture and training procedure and adapt it for each of our domains. BAE-NET has model implementations for 2D binary images and 3D voxel grids, so for these settings we directly use the method as described. For the layout domain we have colored images that can adopt 4 color values (red, green, blue, or grey). In the default version of BAE-NET, it uses an MLP where the second to last layer is size NUM_SEGS and the last layer is size 1; this 1 dimensional output learns a binary occupancy prediction for locations in space. We modify the 2D BAE-NET version so that instead, the second to last layer is still size NUM_SEGS, but the last layer is size 4; in this way we task BAE-NET to solve four binary occupancy problems at once, one for each of our colors. In the layout domain, we still take the part segmentation from BAE-NET by choosing the slot in the second to last layer that activates with the highest potential.
Appendix H Algorithms
We provide pseudo-code references for the key algorithmic components of our method. Algorithm 1 represents our bootstrapped finetuning procedure. Algorithm 2 overviews the inference step of our training loop.
Algorithm 1 Bootstrapped Finetuning
  for outer-loop rounds do
     // Infer template concepts and instantiations
     inf_data = inferConcepts($p_{\text{inf}}$, $X^{*}$)
     // Construct data triplets from inferred data
[…]
     trainGenerative($p_{\text{gen}}$, inf_data)
[…]
     // Finetune inference networks
     for inner-loop rounds do
        batch_data = sampleBatch(
            random.choice([st_data, lest_data, ws_data]))
        MLUpdate($p_{\text{inf}}$, batch_data)
     end for
  end for
Algorithm 2 inferConcepts
  data = []
  for $\widetilde{X}$$\in$ $X^{*}$ do
     $X^{G}$= subSample($\widetilde{X}$, group_size)
     // Search for Template Program that optimizes $O$
[…]
  end for
  return data

### v2

@@ C.1 Omniglot
Representational Capacity The maximum complexity of characters that our method is capable of representing is bounded by (i) the maximum number of tokens that our inference networks can handle and (ii) the maximum number of strokes we sample in the synthetic programs used in the pretraining step. This latter value is set to 12 in our sampling scheme, although through the introduction of HOLE tokens in Template Programs, some of the synthetic programs may end up using more than 12 stroke primitives. While the synthetic pretraining distribution will inform the behavior of the inference networks, this distribution will change over the course of bootstrapped fine-tuning and specialize towards “real” Omniglot examples.
While we observe that these settings allow our model to reliably capture the majority of Omniglot characters, there are some very complex characters that might be hard to fit under these constraints with our top-down inference procedure. It should be possible to relax the constraints of both (i) and (ii), although the cost would be a larger GPU memory footprint and more complex pretraining data, which might require more training time and/or inference networks that use more parameters.
@@
$UBlock\xrightarrow{}\texttt{UNION}(ShBlock,UBlock)\mid ShBlock;$
$ShBlock\xrightarrow{}(SymBlock\mid CBlock\mid MBlock\mid ScBlock);(PBlock\mid UBlock)$
@@
$ScBlock\xrightarrow{}\texttt{Scale}(wt,wf,ht,hf)$
@@
$ShapeBlock\xrightarrow{}(PBlock;ShapeBlock)\mid\texttt{FILL}~{}\mid~{}\texttt{END}$
@@
This DSL creates shape structures by defining cuboids, and arranging them through attachment. Cuboids are instantiated with the Cuboid command. Each Attach command moves one command to connect to previous part, indicated by $cube_{n}$ at a location specified by the other parameters of the command. This language supports the creation of reflectional symmetry groups (Reflect) and translational symmetry groups Translate. Of note, we allow the DSL to expand hierarchically, so that Cuboids can become the bounding volume of their own sub-program (represented above with the return to the START block). These nested sub-programs are allowed to be set to a completely filled mode (FILL) or instead expand into empty space if immediately followed by the END operator. Differing from other languages, we only allow HOLE tokens to replace these START tokens that define sub-program structures, to better match the hierarchical processes by which manufactured shapes are commonly modeled.
@@ D.1 Architecture Details
When we consume a primitive soup of input, we use a different architecture based on a Transformer encoder (Vaswani et al., 2017). We assume that each primitive is a cuboid with 6 dimensions that describe its 3D position and size. We linearize these primitive attributes, and lift each of them to dimension 16 with a 2-layer MLP. Following this we add a learned positional encoding to each attribute based on its attribute type. We then have another ‘positional encoding’ that is produced by concatenating all of the attributes of each primitive (in the lifted dimension) and sending this feature through a 2-layer MLP that outputs an embedding of the same size as the lifted dimension, which then gets summed back into each attribute. This scheme allows us to avoid worrying about how the primitives are ordered, while still allowing the attention scheme of the network to differentiate which attributes belong to which primitives. We send this tokenized representation through a standard Transformer encoder network, where we prepend the sequence with four ‘dummy’ tokens. Each token attends to every other token, and we treat the representations output in the indices of the four ‘dummy’ tokens as the visual tokenization. These dummy tokens build up a representation that attends of the entire input in much the same way as [CLS] tokens have been employed. Note that this encoder assumes a maximum number of primitives as input, which we set to 20. If the input scene does not have 20 primitives, we leave these entries as zeros, and then don’t attend over those corresponding positions in the sequence while encoding.
@@ Appendix E Training Details
We implement our networks in PyTorch (Paszke et al., 2017). We run all experiments on a NVIDIA GeForce RTX 3090 with 24GB of GPU memory, and 64 GB of RAM. During pretraining we set the batch size to max out GPU memory, this amounts to sizes of 32 for the 2D layout domain, 40 for Omniglot domain, 32 for the shape domain with a primitive soup input and 16 for the shape domain with voxel inputs (of size $64^{3}$). Note that this batch size is effectively multiplied by 5 for the ExpansionNet and ParamNet as we train on visual input groups of size 5. During fine-tuning we set the batch size to 20 for all methods, except for the shape-voxels variant, which we set to 10 to avoid maxing out VRAM.
@@
We finetune our inference networks with the procedure described in Section 3.3. For each concept in the training set, we sample a group of visual inputs (at random) from the concept, and record our inference results to produce the LEST and ST dataset. In this way if there are K concepts in $X^{*}$, the size of the ST and LEST data on each training step will also be K. Differing from this, in the wake-sleep step of our finetuning procedure we can generate an arbitrarily large number of paired data by sampling our generative model. We find that sampling a large number of ‘dreams’ is helpful for our finetuning procedure, so we set the number of example $TP$ to sample in each training step to 30,000. This typically takes between 1 and 2 hours, differing slightly for each domain. To encourage the ‘dreams’ we sample to cover a wide-distribution, we design a negative rejection step where we resample any ‘dream’ that either creates an already generated $TP$ or $X^{G}$. We find this rejection criteria is triggered at relatively infrequent rates ($\sim$5of the time).
Once we’ve created the ST, LEST and WS datasets, we use them to finetune our inference networks with cross entropy loss. We train over this datasets for multiple ‘epochs’, where every 5th epoch we run our updated inference networks over concepts from the validation set. We use the Objective $O$ from this validation inference to decide when to break out of the training step, and return to the inference step. This early stopping inference procedure always backtracks to the version of the inference network that achieved the best $O$ measure on the validation set. We use a patience of 10 epochs, and finetune for at most 50 epochs.
Overall, we run our finetuning procedure to convergence for 25, 17, 32 inference-training loops for the layout, omniglot and shape domains respectively. This corresponded with 565, 450, 620 finetuning ‘epochs’ for these domains. For the weights of our objective function $O$, we normalize each reconstruction metric to values typically between 0 and 1, and then we set $\lambda_{1}$ to 1.0 and $\lambda_{2}$ to 0.001. Moreover, when calculating the divergent description length between each Template Program and its respective program instantiations, we discard counting any parameter-types for which we don’t support parameter relations. For instance, as we don’t allow float variables to use parametric relations (see Section B.2), we do not penalize these variables under $O$, because the $TP$ has no opportunity to constrain them.
@@ F.1 Few-shot Generation
Some of these networks directly compare the generated samples to a reference set for each concept. Frechet Distance (FD) (Heusel et al., 2017) measures the distributional similarity between two distributions of encodings. Minimum Matching Distance (MMD) measures the average minimum distance of each member of the reference set to any member of the generated set. Coverage (Cov) measures the percentage of reference set members who are the nearest neighbors to at least one member of the generated set.
We calculate all of the above metrics with respect to a latent space that is domain-specific. To this end, for each domain, we train a visual auto-encoder to learn how to reconstruct ‘random’ scenes from that domain. For the layout domain these are randomly placed primitives. For Omniglot, these are randomly placed strokes. For shapes, these are randomly place cuboids primitives. We train each of these networks to convergence on 500,000 random scenes with a small bottleneck layer size (e.g. 100).
For the layout and omniglot domain we train simple classifier networks to learn a K-way classification over all of the concepts present in the domain. For Omniglot, we train on 19 examples from each of the 1623 characters in the dataset, and hold out one example from each concept as a validation set. Our classifier achieves a 82.4% validation accuracy after convergence. For layout, we train over 95 examples from each concept in a 20-way classification task over meta-concepts; we reach 99.9% validation accuracy on a held out set of 5 examples per concept. The class confidence metric (Conf) is then computed by taking each generated output, running it through the classifier, and then recording the probability that the classifier predicts for the index of the input concept. Note that this metric is not dependant on the reference set of examples.
@@ F.1.1 Perceptual Study
Participants were either shown 50 Omniglot character comparisons or 25 shape comparisons. We visualized shape comparisons with a simple rendering style of the primitive outputs produced by each method (for time considerations).
@@ F.2.1 Ground-Truth Segmentations
We make use of Omniglot stroke data to produce the ground-truth segmentations for characters. We treat each stroke pattern broken by ‘BREAK’ annotations as a separate segment (Lake et al., 2019a). Then, as humans vary in the ways that they order strokes to draw characters, for each test set character we run a clustering procedure to try to find valid and consistent segmentation groupings. We first filter for finding groups of characters that use the same number of strokes, and more than a single stroke (otherwise the co-segmentation task is trivial). Then we encode each stroke with a 4 dimensional feature: its length, its angle, its starting x position, and its starting y position. We run an unsupervised clustering algorithm over this feature representation (Ester et al., 1996), identify if there is any cluster with more than 5 character members, and then take 5 characters from this cluster as a co-segmentation task (where our feature-wise distance creates a correspondence across the strokes of this group). This automatic process generates 306 co-segmentation tasks from the 659 concepts in the Omniglot test set. We manually inspect the generated tasks, and filter out 22 cases where our clustering identified a group that did not have consistent stroke expression. This leaves us with 284 cosegmentation tasks that we use in our experiments.
@@ F.2.2 Group Parsing
Template Programs support parsing by inferring instantiations from a shared $TP$ that explain a group of visual inputs. As each instantiated program $z$ uses the function call structure specified by a Template Program, we can find correspondences in the visual outputs. We create a corresponding group for each primitive type that the Template Program defines: these are created by the PRIM command for the layout domain, the DRAW command for Omniglot and the Cuboid command for 3D Shapes. Note that HOLE tokens are always treated as a construct that creates primitive types. Any command that operates over this primitive type will inherent their corresponding part index (e.g. symmetry operations), excluding combinators like Union.
@@
BAE-NET creates corresponding group parses by performing an argmax over the last layer of an implicit network that is trained to solve occupancy tasks. This implicit network can be run over any spatial position, and assign this input point to one of its part ‘slots’.
@@ G.2 GNS (Feinman & Lake, 2021)
Using the provided code as reference, we re-implemented the system with a simple CNN architecture, following the VAE framing as described in (Rybkin et al., 2020). Though we spent a fair amount of time tuning hyper-parameters, as evidenced by the quantitative results in Table 1, we were unable to achieve competitive performance.
@@
Once we have trained this VQ-VAE for each domain, we can learn our arVHE model. Like the VHE model, and our system, it learns by sampling random visual groups from the same concept. Following the procedure described in the VHE paper and code, we encode these visual concepts with a visual encoder, take a mean embedding, then use this embedding to condition an autoregressive generation process, where the goal is to predict a sequence of VQ-VAE tokens that correspond to another input example from the same concept. We train this network with cross-entropy loss, on the discretized VQ-VAE tokens. For an apples-to-apples comparison against our method, the arVHE baseline uses the same visual encoders that our method uses (Section D). For predicting 3D shapes as a sequence of primitives, we instead just task the VQ-VAE model with predicting discretrized versions of each primitive attribute, where the primitives are randomly ordered (this allows us to skip the VQ-VAE step in this setting).
@@
We follow the author’s implementation released at: https://github.com/czq142857/BAE-NET. We take their architecture and training procedure and adapt it for each of our domains. BAE-NET has model implementations for 2D binary images and 3D voxel grids, so for these settings we directly use the method as described. For the layout domain we have colored images that can adopt 4 color values (red, green, blue, or grey). In the default version of BAE-NET, it uses an MLP where the second to last layer is size NUM_SEGS and the last layer is size 1; this 1 dimensional output learns a binary occupancy prediction for locations in space. We modify the 2D BAE-NET version so that instead, the second to last layer is still size NUM_SEGS, but the last layer is size 4; in this way we task BAE-NET to solve four binary occupancy problems at once, one for each of our colors. In the layout domain, we still take the part segmentation from BAE-NET by choosing the slot in the second to last layer that activates with the highest potential.